Use the Henderson-Hasslebalch equation to calculate the pH of each solution:
(Express your answer using two decimal places)

Part A) A solution that contains 0.620% C5H5N by mass and 0.900% C5H5NHCl by mass (Kb for C5H5N is 1.7 x 10^-9)

Part B) A solution that is 16.5 g of HF and 27.0 g of NaF in 125 mL of solution (Ka for HF is 3.5x10^-4)
Answer #1

A) pKof CH N =-log K = -log(1.7x10 )= 8.77 PK, of C,H-N = 14.00-pky = 14.00 -8.77 = 5.23 - 0.620x10 Concentration of C,H, N,